### v1

# Emotional Images: Assessing Emotions in Images and Potential Biases in Generative Models

Abstract
This paper examines potential biases and inconsistencies in emotional evocation of images produced by generative artificial intelligence (AI) models and their potential bias toward negative emotions. In particular, we assess this bias by comparing the emotions evoked by an AI-produced image to the emotions evoked by prompts used to create those images. As a first step, the study evaluates three approaches for identifying emotions in images—traditional supervised learning, zero-shot learning with vision-language models, and cross-modal auto-captioning—using EmoSet, a large dataset of image-emotion annotations that categorizes images across eight emotional types. Results show fine-tuned models, particularly Google’s Vision Transformer (ViT), significantly outperform zero-shot and caption-based methods in recognizing emotions in images. For a cross-modality comparison, we then analyze the differences between emotions in text prompts—via existing text-based emotion-recognition models—and the emotions evoked in the resulting images. Findings indicate that AI-generated images frequently lean toward negative emotional content, regardless of the original prompt. This emotional skew in generative models could amplify negative affective content in digital spaces, perpetuating its prevalence and impact. The study advocates for a multidisciplinary approach to better align AI emotion recognition with psychological insights and address potential biases in generative AI outputs across digital media.
Introduction
The information environment increasingly relies on visual media, with visual-first platforms like YouTube, Instagram, Pinterest, and TikTok comprising four of the top five most-used social media platforms in the US (Gottfried 2024). At the same time, the academic consensus has increasingly solidified around two findings: 1) the presence of visual media (Casas and Williams 2019; Geise, Panke, and Heck 2021) and 2) the presence of highly emotional content—especially negative—increases engagement in online spaces (Rathje, Bavel, and Linden 2021; Paletz et al. 2023). These findings, taken with the increasing use of social media content and online data as training data for modern large- and visual-language models, have the potential to produce an “unvirtuous cycle” of artificial intelligence (AI) and AI-generated content. In this cycle, content creators produce more emotionally evocative visual media to gain more engagement, generative AI systems then over-represent negative emotions in the media they produce, which then feed back into the information environment, receive more engagement, gain a larger share of visual media we see, and get ingested in the next round of training generative-AI systems.
This paper begins an investigation of this unvirtuous cycle by developing scalable methods to assess emotion evocation in images. To this end, we evaluate three distinct approaches to emotion identification in images—traditional supervised methods, zero-short learning with state-of-the-art vision-language models, and cross-modal auto-captioning approaches—and use these methods to compare emotional salience in textual AI prompts with the emotions evoked in the images produced from these prompts. Leveraging the EmoSet dataset, which contains manually labeled image-emotion pairs for 118,102 images, we demonstrate that fine-tuning computer vision models—particularly Google’s Vision Transformer (ViT) (Dosovitskiy et al. 2021)—substantially outperforms zero-shot and auto-captioning approaches, with textual auto-captioning approaches performing particularly poorly in comparison. Using this fine-tuned ViT model coupled with the state-of-the-art Demux model for emotion classification in text, we compare the emotions present in a text-based generative-AI prompt to the emotions evoked by the produced image. Through this comparison, we test the central hypothesis underlying this unvirtuous cycle: that generative-AI systems are biased toward producing images that evoke negative emotions regardless of the underlying prompt.
Results show that, for the task of emotion recognition in images, fine-tuning off-the-shelf computer vision models, particularly Google’s ViT, substantially outperforms other standard approaches, specifically zero-shot learning using vision-language models and automated captioning approaches. Then, to assess the potential biases between the emotions evoked by AI-generated images and the text prompts that are used to generate them, pulled from Wang et al. (2022), we find, for at least one standard and available generative AI model, the emotions evoked by an image do appear more negative than their source prompts. This result yields evidence of the concerning hypothesis above, namely that a generative AI model may nudge its users and audiences toward negative emotions. This finding is particularly concerning given the state of our modern information environment, where we know negative emotions are contagious (Kramer, Guillory, and Hancock 2014). Hence, we advocate for a multidisciplinary approach to better align AI emotion recognition with psychological insights and address potential biases in generative AI outputs across digital media.
Methods
Datasets for Images, Emotion, and AI Prompts
For model finetuning and initial evaluation, we leverage the EmoSet dataset. Each of the 118,102 images in the multi-class dataset is annotated as one of 8 discreet emotions: Amusement, Awe, Contentment, Excitement, Anger, Disgust, Fear, or Sadness.
To evaluate our central hypothesis that generative AI systems are biased towards negative emotions, we leverage the DiffusionDB dataset: a text-to-image prompt dataset containing 14 million images generated by Stable Diffusion using prompts and hyperparameters specified by real users. We evaluate the salience of emotions within both the prompts and images contained in the 2m-first-5k subset of this dataset.
Emotion Identification in Images
We seek to explore and evaluate different approaches for identifying emotions in AI-generated images, aiming to find the most effective strategies for capturing complex emotional cues. To this end, we leveraged three distinct mechanisms, each targeting a different aspect of emotion recognition. First, we employed zero-shot image classification models (BLIP (Li et al. 2022), CLIP (Radford et al. 2021), and ALBEF) to establish a baseline without task-specific fine-tuning. These models were chosen to assess the potential of large-scale pre-trained models to generalize across a wide variety of emotional contexts. Second, we fine-tuned image classification models such as Google ViT, Microsoft SWIN Transformer, and ConvNeXT (Dosovitskiy et al. 2021; Liu et al. 2021; Woo et al. 2023), focusing on training these models on specialized emotion datasets to enhance their ability to recognize subtle and context-specific emotions. Finally, we implemented an auto-captioning pipeline using BLIP-2 and GPT-4 Vision to generate detailed captions for each image, which were then analyzed by text-based emotion classifiers to evaluate whether adding contextual information could improve emotion recognition accuracy. This presents insights into the curious empirical question of how emotion insight/characteristics are transferred across modalities
To rigorously assess the effectiveness of each approach, we measured macro-average precision, macro-average recall, and macro-average F1 scores. These metrics allowed us to evaluate not only how accurately each model predicted emotions but also how well they generalized across different emotional categories. This three-pronged methodology provided a comprehensive analysis, allowing us to compare the strengths and weaknesses of each approach and gain insights into which strategies were most effective for capturing the diverse range of emotions present in AI-generated images. By combining zero-shot models, fine-tuned classifiers, and a hybrid captioning approach, we aimed to explore the interplay between visual features and textual context in emotion recognition tasks.
Zero-Shot Learning with Vision-Language Models
As an initial step in our study, we employed zero-shot image classification models—BLIP, CLIP, and ALBEF—from the LaViS (Language and Vision) toolkit to establish benchmark results for emotion recognition in AI-generated images. These models were selected for their ability to perform image-text alignment tasks without requiring task-specific fine-tuning, allowing us to rapidly assess the baseline capabilities of general-purpose vision-language models in capturing emotional content.
CLIP (Contrastive Language-Image Pretraining), developed by OpenAI, is trained on a vast dataset of image-text pairs to learn a shared embedding space for images and text. By measuring the similarity between an image and textual descriptions, CLIP can perform zero-shot classification by selecting the label whose text best matches the image. This makes it a powerful tool for tasks requiring semantic understanding without additional training.
ALBEF (Align Before Fuse) enhances image-text alignment through contrastive learning before fusing the modalities for downstream tasks. By aligning visual and textual representations prior to fusion, ALBEF captures nuanced cross-modal relationships, making it effective for tasks like image-text retrieval and, in our case, preliminary emotion recognition.
BLIP (Bootstrapping Language-Image Pre-training) focuses on bootstrapping the learning of vision-language representations using large-scale, noisy web data. It employs a bootstrapping approach that iteratively refines the model’s predictions, improving its understanding of the relationship between images and text. BLIP is versatile for tasks such as image captioning and zero-shot image classification, bridging the gap between vision and language without extensive fine-tuning.
By utilizing these models, we aimed to quickly evaluate the feasibility of extracting emotional content from images using models that leverage large-scale pre-training on diverse datasets. Although these zero-shot models performed less effectively than the fine-tuned image classifiers, their use was instrumental in establishing initial benchmark results. They highlighted the limitations of applying general-purpose models to specialized tasks like emotion recognition, emphasizing the need for models trained on domain-specific data. This initial benchmarking informed our subsequent decision to fine-tune more specialized models.
Fine-Tuning Vision Models for Classification
In this study, we selected Google’s Vision Transformer, ConvNeXT, and Microsoft’s Swin Transformer for fine-tuning on the task of extracting emotions from AI-generated images. These models were chosen due to their diverse architectural designs, strong performance on benchmark datasets, and potential to capture the nuanced features associated with emotional content. By selecting these three models, we aimed to explore a spectrum of architectural paradigms—pure transformers (ViT), modernized CNNs (ConvNeXT), and hybrid architectures (Swin)—to assess their effectiveness in extracting emotions from images.
The Vision Transformer (ViT) applies the transformer architecture to computer vision by treating images as sequences of patches. This allows ViT to leverage self-attention mechanisms to capture global context and long-range dependencies within an image, which is crucial for recognizing emotions that emerge from the overall composition and interrelationships of different regions.
ConvNeXT was selected as a modernized convolutional neural network (CNN) that incorporates design principles from transformers while retaining the efficiency of traditional CNNs. Its convolutional layers excel at capturing local features and fine-grained details, essential for identifying subtle emotional cues conveyed through textures and small visual elements.
The Swin Transformer introduces a hierarchical architecture with shifted windows, enabling it to process images at multiple scales. This multi-scale representation is particularly useful for capturing both coarse and fine details associated with different emotional expressions. Its efficient computation balances modeling power with practical resource usage, making it suitable for complex tasks like emotion recognition.
This diversity in design allows us to evaluate how different approaches handle the emotion extraction task, leveraging their unique strengths. Additionally, all three models have demonstrated state-of-the-art performance on image classification benchmarks, indicating their strong feature extraction capabilities essential for our study. The availability of pre-trained weights for these models facilitates effective fine-tuning on our specialized dataset, which is crucial given the complexity of emotion recognition. Furthermore, the combination of global context modeling (ViT), local feature extraction (ConvNeXT), and multi-scale representation (Swin Transformer) provides a comprehensive toolkit for capturing the multifaceted and subtle nature of emotional content in images.
After fine-tuning, we gather macro-precision, macro-recall, and macro-f1 scores for each base model architecture and select the one with the highest macro f1.
Auto-Captioning Models
To test cross-modality emotion capture, we implement an auto-captioning pipeline to generate textual descriptions of images, which were then analyzed using textual emotion classification models (results evaluated on the DeMuX/MeMO (Chochlakis et al. 2023b, a) model with Semeval (Mohammad et al. 2018), Paletz (Paletz et al. 2024), and GoEmotions (Demszky et al. 2020) models, explained in the following section). We selected BLIP-2 and GPT-4 Vision as our captioning models for this task. The choice of these models was driven by their advanced capabilities in image captioning and their potential to enhance emotion recognition through detailed and context-rich descriptions.
By incorporating BLIP-2 and GPT-4 Vision into our auto-captioning pipeline, we aimed to leverage their respective strengths in generating high-quality image captions to enhance emotion recognition accuracy. BLIP-2’s proven effectiveness in zero-shot tasks and its upgraded capabilities made it a suitable choice for autonomous caption generation. In contrast, GPT-4 Vision’s advanced vision-language integration and the use of a specific prompt allowed us to explore the impact of detailed and exhaustive captions on emotion classification outcomes. This strategic selection enabled us to investigate how the quality and depth of image captions influence the performance of textual emotion classification models, contributing valuable insights to the fields of affective computing and multimodal analysis.
Our decision to include BLIP-2 was influenced by the strong performance of its predecessor, BLIP, in our initial zero-shot model tests for emotion recognition. BLIP demonstrated superior ability among zero-shot models to capture relevant features associated with emotional content in images. Building on this success, we opted to use BLIP-2, the upgraded version, to leverage its improved performance and enhanced capabilities in image captioning.
BLIP-2 excels in generating descriptive and contextually rich captions without the need for supplementary prompts, allowing for an unbiased and autonomous interpretation of the visual content. Its bootstrapping approach iteratively refines predictions, improving the model’s understanding of complex visual cues, including those related to emotions. By utilizing BLIP-2 in our pipeline, we aimed to generate accurate captions that effectively represent the emotional nuances present in the images, providing a solid foundation for subsequent textual emotion classification.
We also selected GPT-4 Vision for its state-of-the-art vision capabilities and proficiency in generating comprehensive and detailed image captions. GPT-4 Vision integrates advanced language understanding with visual processing, enabling it to produce exhaustive descriptions that capture subtle and complex aspects of the image content.
Our rationale for including GPT-4 Vision was to assess whether the exhaustiveness and detail in the captions would significantly impact the accuracy of emotion recognition when analyzed by textual emotion classification models. Given that GPT-4 Vision is at the forefront of current technology, we anticipated that its ability to generate more detailed and nuanced captions would enhance the detection of emotional content by providing richer context and capturing subtle visual cues associated with different emotions.
To focus the captioning process on emotional aspects, we provided GPT-4 Vision with the prompt: ”Generate a descriptive caption for this image that can help a model identify the emotion present.” This directive was intended to guide the model toward highlighting elements relevant to emotion recognition, potentially improving the performance of the subsequent classification models.
Down-Sampling to Minimize GPT Costs
Restricted by API pricing, we opted to only caption 200 images rather than the full validation subset that BLIP2 captioned, allowing us to cost-effectively generate benchmarks.
Emotion Identification in Text
Two points in this research require emotion extraction from text. Firstly, our Auto-captioning pipelines utilize visual image captioning tools to caption a series of images. Those captions are then processed through text-emotion extraction models. Secondly, to extract emotion distributions from the prompts in the DiffusionDB dataset, which we later compare with emotion distributions from the corresponding images.
We leverage three models from the DeMuX-MEmo models to classify text into discrete emotions due to its state-of-the-art ability to handle complex, multi-label emotion recognition tasks. These models were specifically designed to leverage the relationships between emotions, making it an ideal choice for classifying the emotional content present in text. We specifically test models pretrained on the Semeval-2018, Paletz, and GoEmotions benchmarks.
One of the core reasons for selecting DeMuX-MEmo is its superior performance in multi-label settings, where more than one emotion may be present in a single piece of text. Unlike traditional models that treat each label independently, DeMuX-MEmo integrates the correlations between emotions, understanding that certain emotions often co-occur. For instance, emotions like fear and sadness are frequently linked, and recognizing such connections is critical for accurate emotion classification. This capability is particularly valuable in our work, where texts often convey complex emotional states, rather than single, isolated emotions.
DeMuX-MEmo also includes a regularization mechanism that leverages both global and local emotion correlations. By understanding these relationships, the model reduces the risk of making contradictory predictions. For example, if a text is classified as expressing both ”joy” and ”sadness,” the model will use its learned correlations to resolve these conflicting predictions more intelligently than models without this capability. This level of sophistication is essential for our work, where subtle emotional nuances are common.
Comparing Emotion Across Modalities
Assessing whether substantial bias exists in emotions evoked by AI-generated images compared to those that are evoked by the prompts that create them is a key motivation for this work. Directly measuring this divergence is problematic, however, given the different emotions present in the EmoSet dataset versus those in our text-based methods (e.g., the GoEmotion-based pretrained DeMuX-MEmo model includes “joy” and “surprise”, which are not present in the EmoSet labels, and conversely, EmoSet includes “awe”, which is not present in the GoEmotion pretrained dataset). Consequently, we measure two related values: First, we assess Spearman’s $\rho$ rank correlation between emotions in images and the emotions in text, following the validation process outlined in Paletz et al. (2024). In this assessment, we expect positive and significant correlations among related emotions.
As our second evaluation, we directly compare the prevalence of positive and negative emotions present in each modality provided by the Wang et al. (2022) dataset. If our central hypothesis about potential biases toward negative emotions among generative AI systems, the DiffusionDB dataset should over-represent negative emotions in the images compared to text.
Results
Emotion Classification in Images
Table 1 shows results for the emotion-recognition task across our three learning modalities—via multimodal zero-shot models, fine-tuned computer vision models, and automatic image captioning. The zero-shot models in Table 1(a)—BLIP-base, CLIP-ViT-B-16, and ALBEF-base—served as our initial benchmark for emotion recognition without fine-tuning. Across all three models, the overall macro-averaged F1-scores were notably low, indicating limited effectiveness in generalizing emotional cues from the images. Notably, some models demonstrated a tendency to heavily over-predict particular emotions, as evidenced by high class recall but poor class precision for those emotions (emotion recognition performance is provided in the supplemental material). For example, CLIP-ViT-B-16 achieved a recall of 0.722 for “Amusement” but had near-zero recall for “Contentment” and “Disgust”. The model identified a disproportionately high number of images as “Amusement” while struggling to recognize other emotions accurately.
The fine-tuned image classifiers in Table 1(b)—Google’s ViT, Microsoft’s SWIN Transformer, and ConvNeXT—performed significantly better in recognizing emotions than the zero-shot models. These results suggest task-specific training greatly enhanced their ability to capture the diverse range of emotions present in images. Among these models, Google ViT emerged as the strongest performer, with an overall macro-averaged F1-score higher than all other methods.
Lastly, Table 1(c) shows results of the novel use of auto-captioning methods to translate images into text-based captions, where we can leverage the many text-based emotion-recognition models. To assess performance in emotion extraction across the EmoSet dataset, we primarily relied on BLIP-2 (Table 1(c)), as we could run this model on-premises across the full dataset. To assess performance across a variety of text-based emotion-recognition models, we make use of three pre-trained configurations in the Demux-MEmo (Chochlakis et al. 2023b) package, each of which is trained on a separate text-based emotion dataset: GoEmotions (Demszky et al. 2020), Paletz et al. (2024), and SemEval (Mohammad et al. 2018). Regardless of the selected pre-trained model configuration, however, we see that auto-captioning is less performant than the zero-shot multimodal models.
Comparing BLIP2 versus GPT-4o
To investigate potential performance implications from our selection of BLIP2 for automatic captioning, we also generated a sample of image captions using the more expensive and performance GPT-4o interface. Results using GPT-4o are shown in Table 2, where we see that the pre-trained version of Demux-MEmo on the GoEmotions data outperforms the equivalent model using BLIP-2, but the results are not sufficiently higher compared to the zero-shot model to change our approach. This difference in performance may be attributable to more detailed captions compared to BLIP-2. This outcome suggests that the increased descriptiveness of GPT-4o’s captions, which were expected to provide more context and detail for emotion classification, did not lead to a substantial improvement in performance. Looking within specific emotions for instance, the F1-scores for key emotions like “Excitement” and “Fear” were similar across both pipelines. The comparable performance between the BLIP-2 and GPT-4o pipelines indicate that the descriptiveness of captions may not be the primary factor influencing emotion recognition accuracy. Instead, the challenge appears to stem from the inherent difficulty in accurately transferring emotional cues from visual to textual modality.
Model Precision Recall F1
BLIP-base 0.4039 0.3275 0.3150
CLIP-ViT-B16 0.3809 0.2417 0.1695
ALBEF-base 0.3459 0.2371 0.2033
(a) Zero-Shot
Model Precision Recall F1
Google ViT 0.7380 0.7313 0.7343
MS SWIN 0.7065 0.6932 0.6973
ConvNeXT 0.6865 0.6685 0.6740
(b) Fine-Tuned
Model Precision Recall F1
GoEmotions 0.2552 0.2197 0.2154
Paletz 0.1425 0.1808 0.1225
SemEval 0.1525 0.2081 0.1406
(c) Auto-Captioning
Table 1: Image-based Emotion-Recognition Macro-Average Performance Across Learning Strategies (bolded results are highest). Results show fine-tuned computer vision models outperform zero-shot and caption-based methods by a substantial margin when predicting emotion classes in EmoSet (Yang et al. 2023). Auto-captioning methods use BLIP2 to generate captions and Demux-MEmo (Chochlakis et al. 2023b) for text-based emotion extraction.
Model Precision Recall F1
GoEmotions 0.2932 0.2610 0.2236
Paletz 0.1171 0.1531 0.0799
SemEval 0.1224 0.1513 0.0896
Table 2: Sampled Auto-Captioning via GPT-4o. Results are comparable to BLIP-2-based auto-captioning for the GoEmotions pre-trained Demux-MEmo model but are appear substantially below the BLIP-2 for the other pre-trained model configurations.
Comparing Emotions in Prompts and Generated Images
Cross-Correlations Between Modalities
Figure 1 shows the Spearman correlations among pairs of emotions extracted from the prompt (using DeMuX) and from the image (using the Google ViT). This figure suggests negative emotions (anger, disgust, fear, and sadness) align relatively well across these modalities compared to positive emotions. For the positive emotions, joy in prompts corresponds to both amusement and contentment in images, and excited correlates with awe, but surprise and amusement in prompts is less correlated with the positive emotions in images. For context, the scale of correlations seen here—[0.2,+0.2]0.20.2[-0.2,+0.2][ - 0.2 , + 0.2 ]—is close to that in Paletz et al. (2024), where most correlations within the positive and negative sets of emotions are in the 0.3 range.
Figure 1: Cross-Modality Spearman $\rho$ Correlations Between DeMuX and EmoSet. Results show general alignment for the negative emotions (anger, disgust, fear, and sadness). Alignment is less clear in the positive emotions, where joy in prompts corresponds to both amusement and contentment in images, and excited correlates with awe, but surprise and amusement in prompts is less correlated with the positive emotions in images.
For comparison, we also take the highest-performing model by macro-F1 from the Papers with Code leaderboard on text classification in emotions, [^1] https://paperswithcode.com/sota/text-classification-on-emotion the RoBERTa-based Jorgeutd/sagemaker-roberta-base-emotion model. This model uses Amazon’s SageMaker and its Hugging Face Deep Learning container. Results for the negative emotions are quite similar, though the positive emotions do not exhibit as high correlations as with DeMuX (potentially an artifact of a difference in DeMuX’s multi-label approach to discrete emotion classification versus the sagemaker-RoBERTa-base-emotion multi-class construction).
Figure 2: Cross-Modality Spearman $\rho$ Correlations Between Amazon’s SageMaker and EmoSet. Negative emotions appear similarly aligned compared to DeMuX, but positive emotions show less alginment.
Comparing Most Prevalent Emotions
Table 3 compares the rankings of most prevalent emotions in images (via Google ViT) to those present in the prompts. Comparing the emotions evoked by the images to either the DeMuX- or RoBERTa-based models, fear is far more prevalent in the images than in the prompts. Looking specifically at the comparison between Google ViT’s image labels and DeMuX’s prompt labels (where cross correlations are stronger than with the RoBERTa model), fear goes from the most prevalent in the images to the second-to-least prevalent in the prompt. This result yields evidence that, for at least the images from DiffusionDB, AI-generated visuals tend to produce more negative and fearful content than the prompt may otherwise suggest.
Table 3: Cross-Modality Emotion Rankings Between Images and Prompts. Values in the parentheses show average proportions. At least for the DiffusionDB-produced images, those images appear to over-represent fear relative to the emotions evoked by the text-based prompts.
Image Text
Google ViT DeMuX RoBERTa
1 Fear (0.33) Excitement (0.43) Joy (0.54)
2 Amusement (0.20) Joy (0.18) Anger (0.24)
3 Awe (0.12) Surprise (0.11) Fear (0.10)
4 Anger (0.09) Amusement (0.09) Sadness (0.06)
5 Sadness (0.08) Anger (0.06) Surprise (0.04)
6 Excitement (0.07) Disgust (0.05) Love (0.03)
7 Contentment (0.06) Fear (0.05) - -
8 Disgust (0.04) Sadness (0.04) - -